1 he numiti or 10 pts.)Choose the correct answer(Write down the correct answer letter AND value at your answer booklet- For Non-of-the-above choice, write down the correct answer.) 2- 2, ands4 (C) μ,--4 and σ,-16 If randon variable X has a mean μ,-5,astandard deviation σ 2X then: a. vARTHI]EAR R)-(4)(4 (D) r- 6and ơr-3 (E) one of the above elsewhere , then E(X) is 2pexpx- 5/4(E) none of the above 2 If a random variable X has a PDF fx(x)- (B)45 (C 3/5 (D) 5/3 c. For three events A, B, and C, we know that A and C are independent, B and C are independent, A and B are disjoint, P(AUC)-23, P(BUC)-3/4, p(AUBUC)11/ 12, then P(A),P(B), and P(C) are respectively given by:(A)을수 (B),知, 434 43223 2 mutuall y (E) none of the above 322 a exclusi If the continuous random variable X is uniformly distributed with a mean of 45 and a 2p d. variance of 75. The probability that X lies between 40 and 50 is: (E) none of the above You purchase a certain product. The manual states that the lifetime T of the product, defined as the amount of time (in years) the product works properly until it breaks down, satisfies the following rule: P(T 2 t)- e for t 2 0. What is the probability (rounded to nearest thousandth) that this device breaks down during the third year? し(A10.12 2(B)0.54%.-AC) 0.18hi) 0.67(E)noac ofthe above 2nd .--亦 .
Answer #1

(a)

mu_{Y}=E(4-2X)=4-2E(X)=4-2cdot 5=-6

sigma_{Y}=sqrt{Var(4-2X)}=sqrt{4Var(X)}=sqrt{4cdot 4}=4

Correct option is E.

(b)

2x2-2xdx

(c)

From given information we have

P(Acap C)=P(A)P(C),P(Bcap C)=P(B)P(C),P(Acap B)=0

For option D,

3 2 233

2 22 24

12 3 2 2

Hence, correct option is D.

(D)

Let X is uniformly distributed between a and b so

E(X) =-= 45

a+b=90 .........(1)

Var(x) -(a-b

(b - a)2 12 = 75

b-a- 30 ..........(2)

Adding (1) and (2) gives

2b=120

b=60

Putting this in equation (1) gives

а 30

The required probability is

P(40leq xleq 50)=rac{50-40}{60-30}=rac{10}{30}=rac{1}{3}

Correct option is c.

(e)

P(2<X<4)=P(X<4)-P(X<2)=1-P(Xgeq 4)-1+P(Xgeq 2)=e^{-2/5}-e^{-4/5}=0.22

Correct option is E.